The endpoints of the hypotenus of a right triangle are(3,4) and (-2,6). How long is the hypotenuse.
Answer by rfer(10417)

looks like one leg is 2, the other is 5
c^2=2^2+5^2
c^2=4+25
c^2=29
c=5.39